soybean meal is 18% protein cornmeal is 9% how many pounds of each should be mixed together to get 360 lb mixture that is 15% protein
Answer by stanbon(75887)

soybean meal is 18% protein cornmeal is 9% how many pounds of each should be mixed together to get 360 lb mixture that is 15% protein
-----------------------
Equation:
protein + protein = protein
0.18x + 0.09(360-x) = 0.15(360)
--------------------
Multiply thru by 100 to get:
18x + 9*360 - 9x = 15*360
9x = 6*360
---
x = 240 lbs (amt. of soybean meal needed)
-----------------------
360-x = 120 lbs (amt. of cornmeal needed)
===========================================
Cheers,
Stan H.
===========